The claim is that the white blood cell counts of adult females are normally​ distributed, with a standard deviation equal to 3.19. A random sample of 39 adult females has white blood cell counts with a mean of 7.53 and a standard deviation of 3.83. Find the value of the test statistic.

The test statistic is ___

​(Round to three decimal places as​ needed.)
Answer #1

test statistic =\frac{(n-1)s^2}{\sigma^2}

                    =\frac{(39-1)3.19^2}{3.83^2}

                    = 26.361 (ans)